# The $u$-invariant of function fields in one variable

### v1

@@ 3. Valued fields and residue forms
We assume familiarity with standard facts from quadratic form theory over fields as covered in particular by the first chapters of [17] and [10]. On this basis, we will give a succinct introduction to the theory of quadratic forms over valued fields with residue characteristic different from $2$. While in the case of a complete discrete valuation, all these results are well-known, we have not found a presentation of this theory in the literature which would provide the statements in this generality. Our presentation resembles that of [9], a pioneer article on the topic of quadratic forms over valuation rings. The terminology used in [9] is meanwhile a bit off the standard, and the statements in [9] are formulated for a valuation that is assumed to be complete of rank-$1$, while the arguments mostly work in more generality. Our presentation highlights that large parts of the theory may be developed without recourse to a henselization or completion of the valuation ring. To introduce the concepts and formulate a first few facts, we deal with quadratic forms over a local ring in which $2$ is invertible.
Let $R$ be a commutative ring. A quadratic form over $R$ is a homogeneous polynomial $\varphi$ of degree $2$ in a prescribed finite tuple of variables whose length is called the dimension of $\varphi$ and denoted by $\mathsf{dim}(\varphi)$. A quadratic form $\varphi$ over $R$ gives rise to a map $R^{n}\to R$ for $n=\mathsf{dim}(\varphi)$, defined by evaluating $\varphi$ on $n$-tuples, as well as to an $R$-bilinear form
@@
Let $\varphi$ be a quadratic form over $R$. Let $n=\mathsf{dim}(\varphi)$ and $X=(X_{1},\dots,X_{n})$. We say that $\varphi$ represents $a\in R$ if $a=\varphi(x)$ for some $x\in R^{n}$. We call $\varphi$ isotropic if $\varphi(x)=0$ for some $x=(x_{1},\dots,x_{n})\in R^{n}$ such that $R$ is generated as an ideal by the components of $x$. (If $R$ is a local ring, then the latter means that $x_{i}\in{R}^{\times}$ for some $i\in\{1,\dots,n\}$.) If $\varphi$ is not isotropic, then we call $\varphi$ anisotropic.
Let $n\in\mathbb{N}$ and $X=(X_{1},\dots,X_{n})$. Given elements $a_{1},\dots,a_{n}\in R$, we denote the $n$-dimensional quadratic form $a_{1}X_{1}^{2}+\ldots+a_{n}X_{n}^{2}$ over $R$ by $\langle a_{1},\dots,a_{n}\rangle_{R}$, or simply by $\langle a_{1},\dots,a_{n}\rangle$ if the ring $R$ is clear from the context. A quadratic form of this shape (that is, where the terms $X_{i}X_{j}$ with $i\neq j$ all have coefficient 00) is called diagonalized or a diagonal form. Let $\varphi,\psi\in R[X]$ be $n$-dimensional quadratic forms over $R$. We call $\varphi$ and $\psi$ isometric if $\psi(X)=\varphi(X\cdot\alpha)$ for some invertible matrix $\alpha\in\mathbb{M}_{n}(R)$, and we write $\varphi\simeq\psi$ to indicate this. If $\varphi$ and $\psi$ are quadratic forms over $R$, $m=\mathsf{dim}(\varphi)$ and $n=\mathsf{dim}(\psi)$, then we denote by $\varphi\perp\psi$ the $(m+n)$-dimensional quadratic form $\varphi(X_{1},\dots,X_{m})+\psi(X_{m+1},\dots,X_{m+n})$ over $R$. We further denote by $\mathbb{H}_{R}$ or by $\mathbb{H}$ the quadratic form $X_{1}X_{2}$ over $R$.
The next few observations are under the assumption that $2\in{R}^{\times}$. This implies that $\mathbb{H}\simeq_{R}\langle 1,-1\rangle$. Consider an $n$-dimensional quadratic form $\varphi$ in $X=(X_{1},\dots,X_{n})$ over $R$. Then $\varphi=XAX^{\mathsf{t}}$ for a unique symmetric matrix $A\in\mathbb{M}_{n}(R)$. We set $\mathsf{det}(\varphi)=\mathsf{det}(A)$ and call this the determinant of $\varphi$. We call $\varphi$ regular if $A$ is invertible in $\mathbb{M}_{n}(R)$, which is if and only if $\mathsf{det}(A)\in{R}^{\times}$. Note that $\mathsf{det}(\varphi\perp\psi)=\mathsf{det}(\varphi)\cdot\mathsf{det}(\psi)$ for any quadratic form $\psi$ over $R$. Clearly, if $\varphi$ is regular, then so is every quadratic form isometric to $\varphi$.
@@
Clearly, $b$ is represented by $\langle b\rangle_{R}$ and hence also by $\varphi$ if $\varphi\simeq\langle b\rangle_{R}\perp\psi$ for some quadratic form $\psi$ over $R$.
To prove the converse, assume that $\varphi(x)=b$ for some $x=(x_{1},\dots,x_{n})\in R^{n}$. Since $b\in R^{\times}$ and $R$ is local, there exists $i\in\{1,\ldots,n\}$ with $x_{i}\in R^{\times}$. We may without loss of generality assume that $x_{1}\in R^{\times}$.
Let $(e_{1},\ldots,e_{n})$ be the canonical basis of $R^{n}$. We set $e_{1}^{\prime}=x$ and further $e_{i}^{\prime}=e_{i}-(2b)^{-1}\mathfrak{b}_{\varphi}(e_{1}^{\prime},e_{i})e_{1}^{\prime}\in R^{n}$ for $2\leqslant i\leqslant n$. Then $(e_{1}^{\prime},\dots,e_{n}^{\prime})$ is also an $R$-basis of $R^{n}$. Let $\alpha\in\mathbb{M}_{n}(R)$ be the matrix of the base change from $(e_{1}^{\prime},\ldots,e_{n}^{\prime})$ to $(e_{1},\ldots,e_{n})$. Then $\alpha$ is invertible in $\mathbb{M}_{n}(R)$. Using that $\mathfrak{b}_{\varphi}$ is $R$-bilinear, one computes that $\mathfrak{b}_{\varphi}(e_{1}^{\prime},e_{i}^{\prime})=0$ for $2\leqslant i\leqslant n$. Thus, letting $X=(X_{1},\dots,X_{n})$, we have that $\varphi(X)\simeq\varphi(X\alpha)=bX_{1}^{2}+\psi(X_{2},\dots,X_{n})$ for an $(n-1)$-dimensional quadratic form $\psi$ over $R$. ∎
@@
The converse is proven by induction on $n$. For $n=0$, the statement holds trivially. Assume that $n>0$. Let $\mathfrak{m}$ be the maximal ideal of $R$. Then ${R}^{\times}=R\smallsetminus\mathfrak{m}$ and, since $\varphi$ is regular, the entries of the associated matrix cannot all lie in $\mathfrak{m}$. So there exist $x,y\in R^{n}$ with $\mathfrak{b}_{q}(x,y)\notin\mathfrak{m}$. As $\mathfrak{b}_{q}(x,y)=\varphi(x+y)-\varphi(x)-\varphi(y)$, we conclude that $\varphi(z)\notin\mathfrak{m}$ for some $z\in\{x,y,x+y\}$. Letting $a_{1}=\varphi(z)$, we have that $a_{1}\in R\smallsetminus\mathfrak{m}={R}^{\times}$ and obtain by 3.1 that $\varphi\simeq\langle a_{1}\rangle\perp\psi$ for some $(n-1)$-dimensional quadratic form $\psi$ over $R$. As $\varphi$ is regular, so is $\langle a_{1}\rangle\perp\psi$. Hence $a_{1}\mathsf{det}(\psi)\in{R}^{\times}$, which implies that $\mathsf{det}(\psi)\in{R}^{\times}$. Hence $\psi$ is regular. By the induction hypothesis, we obtain that $\psi\simeq\langle a_{2},\dots,a_{n}\rangle_{R}$ for certain $a_{2},\dots,a_{n}\in{R}^{\times}$, whereby $\varphi\simeq\langle a_{1},\dots,a_{n}\rangle_{R}$. ∎
@@ 2. $(b)$
$(a)$ Assume that $\varphi_{K}$ represents $b$. Then $bx_{0}^{2}=a_{1}x_{1}^{2}+\ldots+a_{n}x_{n}^{2}$ for certain $x_{0},\dots,x_{n}\in\mathcal{O}_{v}$, not all contained in $\mathfrak{m}_{v}$. It follows that $x_{i}\in{\mathcal{O}}^{\times}_{v}$ for some $i\in\{1,\dots,n\}$. We obtain by 3.4 that $b=a_{1}y_{1}^{2}+\ldots+a_{n}y_{n}^{2}$ for certain $y_{1},\dots,y_{n}\in\mathcal{O}_{v}$. We conclude by 3.1 that $\varphi\simeq\langle b\rangle_{\mathcal{O}_{v}}\perp\psi$ for some $(n-1)$-dimensional quadratic form $\psi$ over $\mathcal{O}_{v}$. Using 3.2 we obtain that $\psi\simeq\langle b_{2},\dots,b_{n}\rangle_{\mathcal{O}_{v}}$ for certain $b_{2},\dots,b_{n}\in{\mathcal{O}}^{\times}_{v}$. Then $\varphi\simeq\langle b,b_{2},\dots,b_{n}\rangle_{\mathcal{O}_{v}}$.
$(b)$ Assume that $\varphi_{K}$ is isotropic. Hence $a_{1}x_{1}^{2}+\ldots+a_{n}x_{n}^{2}=0$ for certain $x_{1},\dots,x_{n}\in K$ which are not all zero. If $x_{1}=0$, then $\langle a_{2},\dots,a_{n}\rangle_{K}$ is isotropic. As every isotropic form over $K$ represents all elements of $K$, we obtain in any case that $-a_{1}$ is represented by $\langle a_{2},\dots,a_{n}\rangle_{K}$. Applying part $(a)$ to $\langle a_{2},\dots,a_{n}\rangle_{K}$, we obtain that $\langle a_{2},\dots,a_{n}\rangle_{\mathcal{O}_{v}}\simeq\langle-a_{1},b_{3},\dots,b_{n}\rangle_{\mathcal{O}_{v}}$ for certain $b_{3},\dots,b_{n}\in{\mathcal{O}}^{\times}_{v}$. As $\langle a_{1},-a_{1}\rangle_{\mathcal{O}_{v}}\simeq\mathbb{H}_{\mathcal{O}_{v}}$, we conclude that $\varphi\simeq\mathbb{H}_{\mathcal{O}_{v}}\perp\langle b_{3},\dots,b_{n}\rangle_{\mathcal{O}_{v}}$. ∎
@@
A quadratic form over $K$ is $v$-unimodular if and only if its anisotropic part is $v$-unimodular.
@@
An isometry over $\mathcal{O}_{v}$ induces an isometry over $Kv$, so we only need to prove the first part. The proof is by induction on $n$. For $n=0$ there is nothing to prove. Assume that $n\geqslant 1$. As $c_{1}$ is represented by $\langle a_{1},\dots,a_{n}\rangle_{K}$, 3.5 yields that $\langle a_{1},\dots,a_{n}\rangle_{\mathcal{O}_{v}}\simeq\langle c_{1},b_{2},\dots,b_{n}\rangle_{\mathcal{O}_{v}}$ for certain $b_{2},\dots,b_{n}\in{\mathcal{O}}^{\times}_{v}$. Then $\langle c_{1},\dots,c_{n}\rangle_{K}\simeq\langle a_{1},\dots,a_{n}\rangle_{K}\simeq\langle c_{1},b_{2},\dots,b_{n}\rangle_{K}$. Witt’s Cancellation Theorem [17, Chap. I, Theorem 4.2] yields that $\langle c_{2},\dots,c_{n}\rangle_{K}\simeq\langle b_{2},\dots,b_{n}\rangle_{K}$. Now, we obtain from the induction hypothesis, that $\langle c_{2},\dots,c_{n}\rangle_{\mathcal{O}_{v}}\simeq\langle b_{2},\dots,b_{n}\rangle_{\mathcal{O}_{v}}$. Then $\langle a_{1},\dots,a_{n}\rangle_{\mathcal{O}_{v}}\simeq\langle c_{1},b_{2},\dots,b_{n}\rangle_{\mathcal{O}_{v}}\simeq\langle c_{1},c_{2}\dots,c_{n}\rangle_{\mathcal{O}_{v}}$. ∎
@@
Let $\varphi$ be a quadratic form over $K$. Let $\mathcal{C}$ be a set of representatives of ${K}^{\times}/v^{-1}(2vK)$. There exist $r\in\mathbb{N}$, distinct $c_{1},\dots,c_{r}\in\mathcal{C}$, and $v$-unimodular quadratic forms $\varphi_{1},\dots,\varphi_{r}$ over $K$ such that $\varphi\simeq c_{1}\varphi_{1}\perp\ldots\perp c_{r}\varphi_{r}$.
@@
We come to the main statements of this section. They show that, if the valuation $v$ on $K$ is henselian, then unimodular quadratic forms over $K$ are characterised by their residue forms. We will formulate those statements under the hypothesis that $1+\mathfrak{m}_{v}\subseteq{K}^{\times 2}$, which is in particular implied when $v$ is henselian (using that $\mathsf{char}(Kv)\neq 2$); see e.g. [13, Cor. 4.2.4].

### v2

@@ 3. Valued fields and residue forms
We assume familiarity with standard facts from quadratic form theory over fields as covered in particular by the first chapters of [21] and [11]. On this basis, we will give a succinct introduction to the theory of quadratic forms over valued fields with residue characteristic different from $2$. While in the case of a complete discrete valuation all these results are well-known, we have not found a presentation of this theory in the literature which would provide the statements in this generality. Our presentation resembles that of [10], a pioneer article on the topic of quadratic forms over valuation rings. The terminology used in [10] is meanwhile a bit off the standard, and the statements there are formulated for a valuation that is assumed to be complete of rank-$1$, while the arguments mostly work in more generality. Our presentation highlights that large parts of the theory may be developed without recourse to a henselization or completion of the valuation ring. To introduce the concepts and formulate a first few facts, we deal with quadratic forms over a local ring in which $2$ is invertible.
Let $R$ be a commutative ring, with unity. A quadratic form over $R$ is a homogeneous polynomial $\varphi$ of degree $2$ in a prescribed finite tuple of variables whose length is called the dimension of $\varphi$ and denoted by $\mathsf{dim}(\varphi)$. Let $n\in\mathbb{N}$ and consider an $n$-dimensional quadratic form $\varphi$ over $R$. It gives rise to a map $R^{n}\to R$, defined by evaluating $\varphi$ on $n$-tuples, as well as to an $R$-bilinear form
@@
There is a unique symmetric $n\times n$-matrix $A$ over $R$ such that $\mathfrak{b}_{\varphi}$ is given by $(x,y)\mapsto xAy^{\mathsf{t}}$, and we call $\varphi$ nonsingular if $\mathsf{det}(A)\in{R}^{\times}$. We say that $\varphi$ represents $a\in R$ if $a=\varphi(x)$ for some $x\in R^{n}$.
Let $n\in\mathbb{N}$ and $X=(X_{1},\dots,X_{n})$. Given elements $a_{1},\dots,a_{n}\in R$, we denote the $n$-dimensional quadratic form $a_{1}X_{1}^{2}+\ldots+a_{n}X_{n}^{2}$ over $R$ by $\langle a_{1},\dots,a_{n}\rangle_{R}$, or simply by $\langle a_{1},\dots,a_{n}\rangle$ if the ring $R$ is clear from the context. A quadratic form of this shape (that is, where the terms $X_{i}X_{j}$ with $i\neq j$ all have coefficient 0) is called diagonalized or a diagonal form. Let $\varphi,\psi\in R[X]$ be $n$-dimensional quadratic forms over $R$. We call $\varphi$ and $\psi$ isometric if $\psi(X)=\varphi(X\cdot C)$ for some invertible matrix $C\in\mathbb{M}_{n}(R)$, and we write $\varphi\simeq\psi$ to indicate this. If $\varphi$ and $\psi$ are quadratic forms over $R$, $m=\mathsf{dim}(\varphi)$ and $n=\mathsf{dim}(\psi)$, then we denote by $\varphi\perp\psi$ the $(m+n)$-dimensional quadratic form $\varphi(X_{1},\dots,X_{m})+\psi(X_{m+1},\dots,X_{m+n})$ over $R$. We further denote by $\mathbb{H}_{R}$ or by $\mathbb{H}$ the quadratic form $X_{1}X_{2}$ over $R$.
The next few observations are under the assumption that $2\in{R}^{\times}$. Note that this implies that $\mathbb{H}_{R}\simeq\langle 1,-1\rangle$. Consider an $n$-dimensional quadratic form $\varphi$ in $X=(X_{1},\dots,X_{n})$ over $R$. Then $\varphi=XSX^{\mathsf{t}}$ for a unique symmetric matrix $S\in\mathbb{M}_{n}(R)$. We set $\mathsf{det}(\varphi)=\mathsf{det}(S)$ and call this the determinant of $\varphi$. Note that $\mathfrak{b}_{\varphi}:R^{n}\times R^{n}\to R$ is given by $(x,y)\mapsto x(2S)y^{\mathsf{t}}$, and hence $\varphi$ is nonsingular if and only if $\mathsf{det}(S)\in{R}^{\times}$, and we simply call $\varphi$ regular in this case. (We avoid to use the term ‘regular’ without assuming $2\in{R}^{\times}$, as it then occurs with different possible meanings in the literature.) Note that $\mathsf{det}(\varphi\perp\psi)=\mathsf{det}(\varphi)\cdot\mathsf{det}(\psi)$ for any quadratic form $\psi$ over $R$.
@@
Clearly, $b$ is represented by $\langle b\rangle_{R}$ and hence also by $\varphi$ if $\varphi\simeq\langle b\rangle_{R}\perp\psi$ for some quadratic form $\psi$ over $R$. To prove the converse, assume that $\varphi(x)=b$ for some $x=(x_{1},\dots,x_{n})\in R^{n}$. Since $b\in R^{\times}$ and $R$ is local, there exists $i\in\{1,\ldots,n\}$ with $x_{i}\in R^{\times}$. We may without loss of generality assume that $x_{1}\in R^{\times}$. Let $(e_{1},\ldots,e_{n})$ be the canonical basis of $R^{n}$. We set $e_{1}^{\prime}=x$ and further $e_{i}^{\prime}=e_{i}-(2b)^{-1}\mathfrak{b}_{\varphi}(e_{1}^{\prime},e_{i})e_{1}^{\prime}\in R^{n}$ for $2\leqslant i\leqslant n$. Then $(e_{1}^{\prime},\dots,e_{n}^{\prime})$ is also an $R$-basis of $R^{n}$. Let $C\in\mathbb{M}_{n}(R)$ be the matrix of the base change from $(e_{1}^{\prime},\ldots,e_{n}^{\prime})$ to $(e_{1},\ldots,e_{n})$. Then $C$ is invertible in $\mathbb{M}_{n}(R)$. Using that $\mathfrak{b}_{\varphi}$ is $R$-bilinear, one computes that $\mathfrak{b}_{\varphi}(e_{1}^{\prime},e_{i}^{\prime})=0$ for $2\leqslant i\leqslant n$. Thus, letting $X=(X_{1},\dots,X_{n})$, we have that $\varphi(X)\simeq\varphi(X\cdot C)=bX_{1}^{2}+\psi(X_{2},\dots,X_{n})$ for an (n1)(n-1)-dimensional quadratic form $\psi$ over $R$. ∎
@@
The converse is proven by induction on $n$. For $n=0$, the statement holds trivially. Assume that $n>0$. Let $\mathfrak{m}$ be the maximal ideal of $R$. Then ${R}^{\times}=R\smallsetminus\mathfrak{m}$ and, since $\varphi$ is regular, the entries of the associated matrix cannot all lie in $\mathfrak{m}$. So there exist $x,y\in R^{n}$ with $\mathfrak{b}_{q}(x,y)\notin\mathfrak{m}$. As $\mathfrak{b}_{q}(x,y)=\varphi(x+y)-\varphi(x)-\varphi(y)$, we conclude that $\varphi(z)\notin\mathfrak{m}$ for some $z\in\{x,y,x+y\}$. Letting $a_{1}=\varphi(z)$, we have that $a_{1}\in R\smallsetminus\mathfrak{m}={R}^{\times}$ and obtain by 3.1 that $\varphi\simeq\langle a_{1}\rangle\perp\psi$ for some (n1)(n-1)-dimensional quadratic form $\psi$ over $R$. As $\varphi$ is regular, so is $\langle a_{1}\rangle\perp\psi$. Hence $a_{1}\mathsf{det}(\psi)\in{R}^{\times}$, which implies that $\mathsf{det}(\psi)\in{R}^{\times}$. Hence $\psi$ is regular. By the induction hypothesis, we obtain that $\psi\simeq\langle a_{2},\dots,a_{n}\rangle_{R}$ for certain $a_{2},\dots,a_{n}\in{R}^{\times}$, whereby $\varphi\simeq\langle a_{1},\dots,a_{n}\rangle_{R}$. ∎
@@
We collect some standard facts and definitions for quadratic forms over fields, referring to [11, Chap. II] for more details. Let $n\in\mathbb{N}$ and consider an $n$-dimensional quadratic form $\varphi$ over a field $K$. We call $\varphi$ isotropic if there exists $x\in K^{n}\smallsetminus\{0\}$ with $\varphi(x)=0$, anisotropic otherwise. We have that $\varphi$ is nonsingular if and only if it is isometric to $(m\times\mathbb{H}_{K})\perp\psi$ for some natural number $m$ and a nonsingular anisotropic quadratic form $\psi$ over $K$, where $m\times\mathbb{H}_{K}$ denotes the $m$-fold orthogonal sum $\mathbb{H}_{K}\perp\ldots\perp\mathbb{H}_{K}$. This $\psi$ is uniquely determined up to isometry and called the anisotropic part of $\varphi$. A nonsingular quadratic form whose anisotropic part is trivial (0-dimensional) is called hyperbolic. If $\mathsf{char}(K)\neq 2$, then every anisotropic quadratic form over $K$ is nonsingular, and by 3.2 isometric to $\langle a_{1},\ldots,a_{n}\rangle_{K}$ for some $a_{1},\ldots,a_{n}\in{K}^{\times}$.
@@ 2. $(b)$
$(a)$ Assume that $\varphi_{K}$ represents $b$. Then $bx_{0}^{2}=a_{1}x_{1}^{2}+\ldots+a_{n}x_{n}^{2}$ for certain $x_{0},\dots,x_{n}\in\mathcal{O}_{v}$, not all contained in $\mathfrak{m}_{v}$. It follows that $x_{i}\in{\mathcal{O}}^{\times}_{v}$ for some $i\in\{1,\dots,n\}$. We obtain by 3.4 that $b=a_{1}y_{1}^{2}+\ldots+a_{n}y_{n}^{2}$ for certain $y_{1},\dots,y_{n}\in\mathcal{O}_{v}$. We conclude by 3.1 that $\varphi\simeq\langle b\rangle_{\mathcal{O}_{v}}\perp\psi$ for some (n1)(n-1)-dimensional quadratic form $\psi$ over $\mathcal{O}_{v}$. Using 3.2 we obtain that $\psi\simeq\langle b_{2},\dots,b_{n}\rangle_{\mathcal{O}_{v}}$ for certain $b_{2},\dots,b_{n}\in{\mathcal{O}}^{\times}_{v}$. Then $\varphi\simeq\langle b,b_{2},\dots,b_{n}\rangle_{\mathcal{O}_{v}}$.
$(b)$ Assume that $\varphi_{K}$ is isotropic. Hence $a_{1}x_{1}^{2}+\ldots+a_{n}x_{n}^{2}=0$ for certain $x_{1},\dots,x_{n}\in K$ which are not all zero. If $x_{1}=0$, then $\langle a_{2},\dots,a_{n}\rangle_{K}$ is isotropic. As every regular isotropic form over $K$ represents all elements of $K$, we obtain in any case that $a_{1}$ is represented by $\langle a_{2},\dots,a_{n}\rangle_{K}$. Applying part $(a)$ to $\langle a_{2},\dots,a_{n}\rangle_{K}$, we obtain that $\langle a_{2},\dots,a_{n}\rangle_{\mathcal{O}_{v}}\simeq\langle-a_{1},b_{3},\dots,b_{n}\rangle_{\mathcal{O}_{v}}$ for certain $b_{3},\dots,b_{n}\in{\mathcal{O}}^{\times}_{v}$. As $\langle a_{1},-a_{1}\rangle_{\mathcal{O}_{v}}\simeq\mathbb{H}_{\mathcal{O}_{v}}$, we conclude that $\varphi\simeq\mathbb{H}_{\mathcal{O}_{v}}\perp\langle b_{3},\dots,b_{n}\rangle_{\mathcal{O}_{v}}$. ∎
@@
A regular quadratic form over $K$ is $v$-unimodular if and only if its anisotropic part is $v$-unimodular.
@@
An isometry over $\mathcal{O}_{v}$ induces an isometry over $Kv$, so we only need to prove the first part. The proof is by induction on $n$. For $n=0$ there is nothing to prove. Assume that $n\geqslant 1$. As $c_{1}$ is represented by $\langle a_{1},\dots,a_{n}\rangle_{K}$, 3.5 yields that $\langle a_{1},\dots,a_{n}\rangle_{\mathcal{O}_{v}}\simeq\langle c_{1},b_{2},\dots,b_{n}\rangle_{\mathcal{O}_{v}}$ for certain $b_{2},\dots,b_{n}\in{\mathcal{O}}^{\times}_{v}$. Then $\langle c_{1},\dots,c_{n}\rangle_{K}\simeq\langle a_{1},\dots,a_{n}\rangle_{K}\simeq\langle c_{1},b_{2},\dots,b_{n}\rangle_{K}$. Witt’s Cancellation Theorem [21, Chap. I, Theorem 4.2] yields that $\langle c_{2},\dots,c_{n}\rangle_{K}\simeq\langle b_{2},\dots,b_{n}\rangle_{K}$. Now, we obtain from the induction hypothesis, that $\langle c_{2},\dots,c_{n}\rangle_{\mathcal{O}_{v}}\simeq\langle b_{2},\dots,b_{n}\rangle_{\mathcal{O}_{v}}$. Then $\langle a_{1},\dots,a_{n}\rangle_{\mathcal{O}_{v}}\simeq\langle c_{1},b_{2},\dots,b_{n}\rangle_{\mathcal{O}_{v}}\simeq\langle c_{1},c_{2}\dots,c_{n}\rangle_{\mathcal{O}_{v}}$. ∎
@@
Let $\varphi$ be a regular quadratic form over $K$. Let $\mathcal{C}$ be a set of representatives of ${K}^{\times}/v^{-1}(2vK)$. There exist $r\in\mathbb{N}$, distinct $c_{1},\dots,c_{r}\in\mathcal{C}$, and $v$-unimodular quadratic forms $\varphi_{1},\dots,\varphi_{r}$ over $K$ such that $\varphi\simeq c_{1}\varphi_{1}\perp\ldots\perp c_{r}\varphi_{r}$.
@@
We come to the main statements of this section. They show that, if the valuation $v$ on $K$ is henselian, then unimodular quadratic forms over $K$ are characterised by their residue forms. We will formulate those statements under the hypothesis that $1+\mathfrak{m}_{v}\subseteq{K}^{\times 2}$. This hypothesis holds in particular when $v$ is henselian. (It is equivalent to $v$ being $2$-henselian; see 7.3.)